# Learning with Calibration: Exploring Test-Time Computing of Spatio-Temporal Forecasting

### v1

@@ D.1 Datasets Details
Our experiments are carried out on 14 real-world datasets from diffrent domain. The statistics of these spatio-temporal datasets are shown in Table 5.
Table 5: Summary of datasets used for our experiments. Degree: the average degree of each node. Meta: the number of metadata associated with each node. Data Points: multiplication of nodes and frames. M: million ($10^{6}$).
@@
 [37] UrbanEV 275 09/01/2022 – 02/28/2023 4344 1 hour 1.19M
 [77] Know-Air 184 01/01/2015 – 12/31/2018 11688 3 hours 2.15M
 [38] METR-LA 207 03/01/2012 – 06/27/2012 34,272 5 minutes 7.09M
LargeST [46] CA 8,600 01/01/2019 – 12/31/2019 35,040 15 minutes 30.13M
@@
 [8] Air-Stream 1087 $\rightarrow$ 1154 $\rightarrow$ 1193 $\rightarrow$ 1202 01/01/2016 - 12/31/2019 34065 1 hour 15.79M
PEMS-Stream 655 $\rightarrow$ 715 $\rightarrow$ 786 $\rightarrow$ 822 $\rightarrow$ 834 $\rightarrow$ 850 $\rightarrow$ 871 07/10/2011 - 09/08/2017 61,992 5 minutes 34.30M
Energy-Stream 103 $\rightarrow$ 113 $\rightarrow$ 122 $\rightarrow$ 134 Unknown (245 days) 34,560 10 minutes 1.63M
We follow the conventional practice [38] to define the graph topology for all spatio-temporal datasets except Know-Air. Specifically, we construct the adjacency matrix $A$ for each dataset using a threshold Gaussian kernel, defined as follows:
$A_{[ij]}=\begin{cases}\exp\left(-\frac{d_{ij}^{2}}{\sigma^{2}}\right)&\text{if}\exp\left(-\frac{d_{ij}^{2}}{\sigma^{2}}\right)\geq r~{}\text{and}~{}i\neq j\\ 0&\text{otherwise}\end{cases}$
For the KnowAir dataset, we follow the original paper [77] and calculate the correlation between nodes to construct the adjacency matrix. Intuitively, most aerosol pollutants are distributed within a certain range above the ground. In addition, the mountains along the two cities will hinder the transmission of pollutants to the PM2.5 direction. Based on these intuitions, we constrain the weights in the adjacency matrix by the following formula:
$d_{ij}=||\rho_{i}-\rho_{j}||,\quad m_{ij}=\sup_{\lambda\in(0,1)}\{h(\lambda\rho_{i}+(1-\lambda)\rho_{j})-\max\{h(\rho_{i}),h(\rho_{j})\}\},$
where $\rho_{i}$ is the location (latitude, longitude) of node $i$, $h(\rho)$ is the height of location $\rho$, and |||||$\cdot$||| | ⋅ | | is the L2-norm of the vector. $H(\cdot)$ is the Heaviside step function, where $H(x)=1$ if and only if $x>0$. $d_{\theta}$ and $m_{\theta}$ are the distance and altitude thresholds, respectively. Specifically, we also set the distance threshold $d_{\theta}=300$ km and the altitude threshold $m_{\theta}=1200$ meters.
@@ D.2 Baseline Details
  • STAEformer [45]: STAEformer is a spatial-temporal adaptive embedding transformer that makes vanilla transformer state-of-the-art for spatio-temporal forecasting. It introduces a novel architecture to effectively capture the dynamic spatial and temporal dependencies in spatio-temporal data. https://github.com/XDZhelheim/STAEformer
  • STTN [83]: STTN is a spatial-temporal transformer network designed for traffic flow forecasting. It leverages dynamic directed spatial dependencies and long-range temporal dependencies to enhance the accuracy of long-term traffic predictions. https://github.com/xumingxingsjtu/STTN
  • GWNet [80]: GWNet is a graph wavenet model for deep spatial-temporal graph modeling. It effectively captures the complex spatial and temporal patterns in spatio-temporal data using a combination of graph convolutional networks and dilated causal convolutions. https://github.com/nnzhan/Graph-WaveNet
  • STGCN [87]: STGCN is a spatio-temporal graph convolutional network framework for traffic forecasting. It integrates graph convolutional networks with temporal convolutional networks to model the spatial and temporal dependencies in traffic data. https://github.com/hazdzz/stgcn
  • STID [60]: STID is a simple yet effective baseline for spatio-temporal forecasting. It addresses the indistinguishability of samples in spatial and temporal dimensions by attaching spatial and temporal identity information, achieving competitive performance with concise and efficient models. https://github.com/GestaltCogTeam/STID
  • ST-Norm [13]: ST-Norm is a method that applies spatial and temporal normalization for multi-variate time series forecasting. It enhances the performance of forecasting models by normalizing the spatial and temporal features of the data. https://github.com/JLDeng/ST-Norm
@@
  • PatchSTG [15]: PatchSTG is an attention-based dynamic spatial modeling method that uses irregular spatial patching for efficient large-scale spatio-temporal forecasting. It reduces computational complexity by segmenting large-scale inputs into balanced and non-overlapped patches, capturing local and global spatial dependencies effectively. https://github.com/lmissher/patchstg
@@
  • STONE [66]: STONE is a state-of-the-art spatio-temporal OOD learning framework that effectively models spatial heterogeneity and generates temporal and spatial semantic graphs. It introduces a graph perturbation mechanism to enhance the model’s environmental modeling capability for better generalization. https://github.com/PoorOtterBob/STONE-KDD-2024
@@
  • EAC [8]: EAC is a state-of-the-art method for exploring the rapid adaptation of models in the face of dynamic spatio-temporal graph changes during supervised finetuning. It follows the principles of expand and compress to address the challenges of retraining models over new data and catastrophic forgetting. https://github.com/Onedean/EAC
  • STKEC [68]: STKEC is a continual learning framework for traffic flow prediction on expanding traffic networks. It introduces a pattern bank to store representative network patterns and employs a pattern expansion mechanism to incorporate new patterns from evolving networks without requiring historical graph data. https://github.com/wangbinwu13116175205/STKEC
@@
  • TTT-MAE [18]: TTT-MAE is a test-time training method that uses masked autoencoders to adjust the model during inference. It helps improve the performance of the model on unseen data by effectively utilizing test-time information. We adapt it to the backbone model of the spatiotemporal network, which is divided into a feature extractor and a prediction head as well as a self-supervisory head. https://github.com/Rima-ag/TTT-MAE
  • TENT [71]: TENT is a method for adjusting the model at test time by normalizing the activation function to reduce the offset between the training distribution and the test distribution. It enhances the generalization ability of the model without retraining on labeled test data. Although it is theoretically designed mainly for the cross entropy loss function, that is, classification tasks, we can still directly apply it to our prediction scenarios. https://github.com/DequanWang/tent
@@
  • OnlineTCN [101]: OnlineTCN is an online learning method based on a time convolutional network. It can adapt to new data sequentially and is very suitable for real-time prediction applications where data arrives continuously. https://github.com/locuslab/TCN
  • FSNet [55]: FSNet proposes a fast and slow learning network for online time series prediction that can handle both sudden changes and repeated patterns. In particular, FSNet improves on a slowly learning backbone by dynamically balancing fast adaptation to recent changes and retrieval of similar old knowledge. FSNet implements this mechanism through the interaction between two complementary components of the adapter to monitor each layer’s contribution to missing events, and an associative memory that supports remembering, updating, and recalling repeated events. https://github.com/salesforce/fsnet
  • OneNet [78]: OneNet dynamically updates and combines two models, one focusing on modeling dependencies across time dimensions and the other focusing on cross-variable dependencies. The approach integrates reinforcement learning-based methods into a traditional online convex programming framework, allowing the two models to be linearly combined with dynamically adjusted weights, thereby addressing the main drawback of classical online prediction methods that are slow to adapt to concept drift. https://github.com/yfzhang114/OneNet
@@
  • CompFormer [97]: CompFormer proposes a test-time compensated representation learning framework, including a spatiotemporal decomposed database and a multi-head spatial transformer model. The former component explicitly separates all training data along the time dimension according to periodic features, while the latter component establishes connections between recent observations and historical sequences in the database through a spatial attention matrix. This enables it to transfer robust features to overcome abnormal events
  • DOST [70]: DOST proposes a novel online continuous learning framework tailored to the characteristics of spatiotemporal data. DOSTadopts an adaptive spatiotemporal network equipped with variable independent adapters to dynamically address the unique distribution changes of each urban location. In addition, to adapt to the gradual nature of these transformations, a wake-sleep learning strategy is used, which intermittently fine-tunes the adapters during the online stage to reduce computational overhead.
@@ D.3 Protocol Details
$\text{MAE}=\frac{1}{n}\sum_{i=1}^{n}|y_{i}-\hat{y}_{i}|,~{}~{}~{}\text{RMSE}=\sqrt{\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}},~{}~{}~{}\text{MAPE}=\frac{100\%}{n}\sum_{i=1}^{n}\left|\frac{\hat{y}_{i}-y_{i}}{y_{i}}\right|$
@@
Parameter Detail. For the hyper-parameter settings of all baseline methods, we follow the parameter settings recommended by the corresponding references. For our paper, except for the robustness study section, all other experimental hyper-parameters are set uniformly: the learning rate $lr$ is 1e-4, and the number of groups $m$ is set to 4. All experiments are conducted on a Linux server equipped with a 1 × AMD EPYC 7763 128-Core Processor CPU (256GB memory) and 4 × NVIDIA RTX A6000 (48GB memory) GPUs. To carry out benchmark testing experiments, all baselines are set to run for a duration of 100$\sim$150 epochs by default (depends on the corresponding paper), with specific timings contingent upon the method with early stop mechanism. The number of early stopping steps is set to 10.
@@ E.1 Complete Results Table
We provide complete information of the experimental tables in the main text as Table LABEL:tab:rq2_fewshot_1278
Table 6: Performance comparison of different models w/ and w/o ST-TTC in the few-shot scenario.
@@
STAEformer [45] STTN [83] GWNet [80] STGCN [87] STID [60] ST-Norm [13]
@@
PEMS-03 MAE $23.57{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.90}}$ 22.96±0.93 $21.32{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.93}}$ 21.15±0.92 $21.70{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.98}}$ 21.43±0.92 $21.79{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.50}}$ 21.28±0.54 $21.81{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.24}}$ 21.57±0.24 $21.13{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.31}}$ 20.74±0.28
@@
RMSE $37.90{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.41}}$ 37.10±1.59 $34.19{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.59}}$ 33.87±1.53 $34.52{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.44}}$ 34.28±1.40 $34.82{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.92}}$ 34.08±0.89 $35.58{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.53}}$ 35.18±0.52 $33.76{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.45}}$ 33.07±0.35
@@
MAPE(%) $21.49{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.21}}$ 21.23±1.03 $21.14{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.35}}$ 21.13±1.28 $21.30{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.14}}$ 19.70±0.57 $22.85{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.76}}$ 21.59±0.77 $21.46{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.86}}$ 21.27±0.73 $22.57{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±3.13}}$ 22.04±2.22
@@
PEMS-04 MAE $35.10{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±4.25}}$ 34.57±4.08 $29.76{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.37}}$ 29.44±0.31 $33.22{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±1.86}}$ 32.87±1.95 $29.97{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.81}}$ 29.66±0.83 $29.64{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.67}}$ 29.49±0.65 $30.66{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{±0.11}}$ 30.31±0.14

### v2

@@ D.1 Datasets Details
Our experiments are carried out on 14 real-world datasets from diffrent domain. The statistics of these spatio-temporal datasets are shown in Table D.1.
We follow the conventional practice (Li et al., 2018) to define the graph topology for all spatio-temporal datasets except Know-Air. Specifically, we construct the adjacency matrix $A$ for each dataset using a threshold Gaussian kernel, defined as follows:
$A_{[ij]}=\begin{cases}\exp\left(-\frac{d_{ij}^{2}}{\sigma^{2}}\right)&\text{if }\exp\left(-\frac{d_{ij}^{2}}{\sigma^{2}}\right)\geq r\penalty 10000\ \text{and}\penalty 10000\ i\neq j\\ 0&\text{otherwise}\end{cases}$
For the KnowAir dataset, we follow the original paper (Wang et al., 2020) and calculate the correlation between nodes to construct the adjacency matrix. Intuitively, most aerosol pollutants are distributed within a certain range above the ground. In addition, the mountains along the two cities will hinder the transmission of pollutants to the PM2.5 direction. Based on these intuitions, we constrain the weights in the adjacency matrix by the following formula:
[…]
where $\rho_{i}$ is the location (latitude, longitude) of node $i$, $h(\rho)$ is the height of location $\rho$, and $||\cdot||$ is the L2-norm of the vector. $H(\cdot)$ is the Heaviside step function, where $H(x)=1$ if and only if $x>0$. $d_{\theta}$ and $m_{\theta}$ are the distance and altitude thresholds, respectively. Specifically, we also set the distance threshold $d_{\theta}=300$ km and the altitude threshold $m_{\theta}=1200$ meters.
Table D.1: Summary of datasets used for our experiments. Degree: the average degree of each node. Meta: the number of metadata associated with each node. Data Points: multiplication of nodes and frames. M: million ($10^{6}$).
@@
(Li et al., 2025) UrbanEV 275 09/01/2022 – 02/28/2023 4344 1 hour 1.19M
(Wang et al., 2020) Know-Air 184 01/01/2015 – 12/31/2018 11688 3 hours 2.15M
(Li et al., 2018) METR-LA 207 03/01/2012 – 06/27/2012 34,272 5 minutes 7.09M
LargeST (Liu et al., 2023b) CA 8,600 01/01/2019 – 12/31/2019 35,040 15 minutes 30.13M
@@
(Chen and Liang, 2025) Air-Stream
1087 $\rightarrow$ 1154
$\rightarrow$ 1193 $\rightarrow$ 1202
01/01/2016 - 12/31/2019 34065 1 hour 15.79M
PEMS-Stream
655 $\rightarrow$ 715 $\rightarrow$ 786
$\rightarrow$ 822 $\rightarrow$ 834 $\rightarrow$ 850
$\rightarrow$ 871
07/10/2011 - 09/08/2017 61,992 5 minutes 34.30M
Energy-Stream
103 $\rightarrow$ 113
$\rightarrow$ 122 $\rightarrow$ 134
Unknown (245 days) 34,560 10 minutes 1.63M
@@ D.2 Baseline Details
  • STAEformer (Liu et al., 2023a): STAEformer is a spatial-temporal adaptive embedding transformer that makes vanilla transformer state-of-the-art for spatio-temporal forecasting. It introduces a novel architecture to effectively capture the dynamic spatial and temporal dependencies in spatio-temporal data. https://github.com/XDZhelheim/STAEformer
  • STTN (Xu et al., 2020): STTN is a spatial-temporal transformer network designed for traffic flow forecasting. It leverages dynamic directed spatial dependencies and long-range temporal dependencies to enhance the accuracy of long-term traffic predictions. https://github.com/xumingxingsjtu/STTN
  • GWNet (Wu et al., 2019): GWNet is a graph wavenet model for deep spatial-temporal graph modeling. It effectively captures the complex spatial and temporal patterns in spatio-temporal data using a combination of graph convolutional networks and dilated causal convolutions. https://github.com/nnzhan/Graph-WaveNet
  • STGCN (Yu et al., 2018): STGCN is a spatio-temporal graph convolutional network framework for traffic forecasting. It integrates graph convolutional networks with temporal convolutional networks to model the spatial and temporal dependencies in traffic data. https://github.com/hazdzz/stgcn
  • STID (Shao et al., 2022b): STID is a simple yet effective baseline for spatio-temporal forecasting. It addresses the indistinguishability of samples in spatial and temporal dimensions by attaching spatial and temporal identity information, achieving competitive performance with concise and efficient models. https://github.com/GestaltCogTeam/STID
  • ST-Norm (Deng et al., 2021): ST-Norm is a method that applies spatial and temporal normalization for multi-variate time series forecasting. It enhances the performance of forecasting models by normalizing the spatial and temporal features of the data. https://github.com/JLDeng/ST-Norm
@@
  • PatchSTG (Fang et al., 2025): PatchSTG is an attention-based dynamic spatial modeling method that uses irregular spatial patching for efficient large-scale spatio-temporal forecasting. It reduces computational complexity by segmenting large-scale inputs into balanced and non-overlapped patches, capturing local and global spatial dependencies effectively. https://github.com/lmissher/patchstg
@@
  • STONE (Wang et al., 2024a): STONE is a state-of-the-art spatio-temporal OOD learning framework that effectively models spatial heterogeneity and generates temporal and spatial semantic graphs. It introduces a graph perturbation mechanism to enhance the model’s environmental modeling capability for better generalization. https://github.com/PoorOtterBob/STONE-KDD-2024
@@
  • EAC (Chen and Liang, 2025): EAC is a state-of-the-art method for exploring the rapid adaptation of models in the face of dynamic spatio-temporal graph changes during supervised finetuning. It follows the principles of expand and compress to address the challenges of retraining models over new data and catastrophic forgetting. https://github.com/Onedean/EAC
  • STKEC (Wang et al., 2023a): STKEC is a continual learning framework for traffic flow prediction on expanding traffic networks. It introduces a pattern bank to store representative network patterns and employs a pattern expansion mechanism to incorporate new patterns from evolving networks without requiring historical graph data. https://github.com/wangbinwu13116175205/STKEC
@@
  • TTT-MAE (Gandelsman et al., 2022): TTT-MAE is a test-time training method that uses masked autoencoders to adjust the model during inference. It helps improve the performance of the model on unseen data by effectively utilizing test-time information. We adapt it to the backbone model of the spatiotemporal network, which is divided into a feature extractor and a prediction head as well as a self-supervisory head. https://github.com/Rima-ag/TTT-MAE
  • TENT (Wang et al., 2021): TENT is a method for adjusting the model at test time by normalizing the activation function to reduce the offset between the training distribution and the test distribution. It enhances the generalization ability of the model without retraining on labeled test data. Although it is theoretically designed mainly for the cross entropy loss function, that is, classification tasks, we can still directly apply it to our prediction scenarios. https://github.com/DequanWang/tent
@@
  • OnlineTCN (Zinkevich, 2003): OnlineTCN is an online learning method based on a time convolutional network. It can adapt to new data sequentially and is very suitable for real-time prediction applications where data arrives continuously. https://github.com/locuslab/TCN
  • FSNet (Pham et al., 2023): FSNet proposes a fast and slow learning network for online time series prediction that can handle both sudden changes and repeated patterns. In particular, FSNet improves on a slowly learning backbone by dynamically balancing fast adaptation to recent changes and retrieval of similar old knowledge. FSNet implements this mechanism through the interaction between two complementary components of the adapter to monitor each layer’s contribution to missing events, and an associative memory that supports remembering, updating, and recalling repeated events. https://github.com/salesforce/fsnet
  • OneNet (Wen et al., 2023): OneNet dynamically updates and combines two models, one focusing on modeling dependencies across time dimensions and the other focusing on cross-variable dependencies. The approach integrates reinforcement learning-based methods into a traditional online convex programming framework, allowing the two models to be linearly combined with dynamically adjusted weights, thereby addressing the main drawback of classical online prediction methods that are slow to adapt to concept drift. https://github.com/yfzhang114/OneNet
@@
  • CompFormer (Zhang et al., 2023c): CompFormer proposes a test-time compensated representation learning framework, including a spatiotemporal decomposed database and a multi-head spatial transformer model. The former component explicitly separates all training data along the time dimension according to periodic features, while the latter component establishes connections between recent observations and historical sequences in the database through a spatial attention matrix. This enables it to transfer robust features to overcome abnormal events
  • DOST (Wang et al., 2024c): DOST proposes a novel online continuous learning framework tailored to the characteristics of spatiotemporal data. DOSTadopts an adaptive spatiotemporal network equipped with variable independent adapters to dynamically address the unique distribution changes of each urban location. In addition, to adapt to the gradual nature of these transformations, a wake-sleep learning strategy is used, which intermittently fine-tunes the adapters during the online stage to reduce computational overhead.
@@ D.3 Protocol Details
$\text{MAE}=\frac{1}{n}\sum_{i=1}^{n}|y_{i}-\hat{y}_{i}|,\penalty 10000\ \penalty 10000\ \penalty 10000\ \text{RMSE}=\sqrt{\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}},\penalty 10000\ \penalty 10000\ \penalty 10000\ \text{MAPE}=\frac{100\%}{n}\sum_{i=1}^{n}\left|\frac{\hat{y}_{i}-y_{i}}{y_{i}}\right|$
@@
Parameter Detail. For the hyper-parameter settings of all baseline methods, we follow the parameter settings recommended by the corresponding references. For our paper, except for the robustness study section, all other experimental hyper-parameters are set uniformly: the learning rate $lr$ is 1e-4, and the number of groups $m$ is set to 4. All experiments are conducted on a Linux server equipped with a 1 × AMD EPYC 7763 128-Core Processor CPU (256GB memory) and 4 × NVIDIA RTX A6000 (48GB memory) GPUs. To carry out benchmark testing experiments, all baselines are set to run for a duration of $0\sim1$50 epochs by default (depends on the corresponding paper), with specific timings contingent upon the method with early stop mechanism. The number of early stopping steps is set to 10.
@@ E.1 Complete Results Table
We provide complete information of the experimental tables in the main text as Table LABEL:tab:rq2_fewshot_12, E.2, E.3
Table E.1: Performance comparison of different models w/ and w/o ST-TTC in the few-shot scenario.
@@
STAEformer (Liu et al., 2023a) STTN (Xu et al., 2020) GWNet (Wu et al., 2019) STGCN (Yu et al., 2018) STID (Shao et al., 2022b) ST-Norm (Deng et al., 2021)
@@
PEMS-03 MAE $3.57{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.90}} 22$.96±0.93 $21.32{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.93}}$ 21.15±0.92 $21.70{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.98}}$ 21.43±0.92 $21.79{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.50}}$ 21.28±0.54 $21.81{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.24}}$ 21.57±0.24 $21.13{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.31}}$ 20.74±0.28
@@
RMSE $37.90{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 1.41}}$ 37.10±1.59 $4.19{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 1.59}} 33$.87±1.53 $34.52{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 1.44}}$ 34.28±1.40 $34.82{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.92}}$ 34.08±0.89 $35.58{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.53}}$ 35.18±0.52 $3.76{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.45}} 33$.07±0.35
@@
MAPE(%) $21.49{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 1.21}}$ 21.23±1.03 $21.14{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 1.35}}$ 21.13±1.28 $21.30{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 1.14}}$ 19.70±0.57 $22.85{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.76}}$ 21.59±0.77 $21.46{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.86}}$ 21.27±0.73 $2.57{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 3.13}} 22$.04±2.22
@@
PEMS-04 MAE $35.10{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 4.25}}$ 34.57±4.08 $29.76{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.37}}$ 29.44±0.31 $33.22{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 1.86}}$ 32.87±1.95 $29.97{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.81}}$ 29.66±0.83 $29.64{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.67}}$ 29.49±0.65 $30.66{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{\scriptsize\textpm 0.11}}$ 30.31±0.14